# Beyond Correlation: Causal Multi-View Unsupervised Feature Selection Learning

### v1

@@ 5.2 Experiments on Synthetic Dataset
Following common data generation protocols in causal literature [46, 47], we construct a synthetic multi-view dataset to evaluate the effectiveness of CAUSA in mitigating spurious correlations and identifying causal features. The synthetic dataset consists of 500 samples with two views, each view containing 10 causal features. In addition, the first and second views contain 410 and 390 non-causal features, respectively, of which 260 and 240 are spurious features and 150 are noise features. The causal features of each view are independently sampled from Gaussian distributions and directly determine the generation of class labels, thereby establishing causal dependencies between the features and the labels. Spurious features are generated from confounding factors, yielding apparent but non-causal correlations with the labels. Noise features are randomly sampled from a Gaussian distribution and have negligible correlations with the labels. Fig. 3 illustrates the feature selection results of CAUSA and the runner-up method on the synthetic dataset. The experimental results indicate that CAUSA effectively identifies causal features, while the runner-up method frequently selects non-causal features with spurious correlations.

### v2

@@ 5.2 Experiments on Synthetic Dataset
Following common data generation protocols in causal literature [lu2021invariant, kallus2018causal], we construct a synthetic multi-view dataset to evaluate the effectiveness of CAUSA in mitigating spurious correlations and identifying causal features. The synthetic dataset consists of 500 samples with two views, each view containing 10 causal features. In addition, the first and second views contain 410 and 390 non-causal features, respectively, of which 260 and 240 are spurious features and 150 are noise features. The causal features of each view are independently sampled from Gaussian distributions and directly determine the generation of class labels, thereby establishing causal dependencies between the features and the labels. Spurious features are generated from confounding factors, yielding apparent but non-causal correlations with the labels. Noise features are randomly sampled from a Gaussian distribution and have negligible correlations with the labels. Fig. 3 illustrates the feature selection results of CAUSA and the runner-up method on the synthetic dataset. The experimental results indicate that CAUSA effectively identifies causal features, while the runner-up method frequently selects non-causal features with spurious correlations.